OpenStudy (anonymous):

Lim (y-π/2)square / 1-siny x->π/2 How can I get the limit?

OpenStudy (anonymous):

@Coke How come the 'x' approaches pi/2 but the variable in the function is 'y'?

OpenStudy (anonymous):

sorry its y approaches pi/2

OpenStudy (dumbcow):

hope this is calculus class....L'hopital rule :)

OpenStudy (anonymous):

yes sir its calculus

OpenStudy (theeric):

\[\Large\lim\limits_{y\rightarrow \frac{\pi}{2}}=\frac{\left(y-\frac{\pi}{2}\right)^2}{1-\sin y}\] Thought I'd make it "pretty." :P

OpenStudy (dumbcow):

differentiate top/bottom separately then reapply limit

OpenStudy (anonymous):

thats it sir thanks :D

OpenStudy (anonymous):

\[\bf \lim_{y \rightarrow \frac{\pi}{2}}\frac{ (y-\frac{\pi}{2})^2 }{1-\sin(y) }\]Use l'hopital's and differentiate the top and bottom twice since you get 0/0 after differentiating once:\[ \bf =\lim_{y \rightarrow \frac{\pi}{2}}\frac{ 2 }{ \sin(y) }=\frac{ 2 }{ \sin(\pi/2) }=2\]

OpenStudy (anonymous):

love the way he keeps saying "sir"...lol @Coke By any chance, is @dumbcow your calc professor? =P

OpenStudy (anonymous):

how did you get the 2 in the numerator sir I cant get it

OpenStudy (anonymous):